You vaporize a liquid substance at 100 oC and 755 mmHg. The volume of 0.548 g of vapor is 237 mL. What is the molecular mass of the substance?
Answer #1

T = 100 + 273 = 373 K

P = 755 mmH = 755 /760 = 0.9934 atm

V = 237 mL = 0.237 L

R = 0.0821 L atm / mol K

ideal gas equation :

P V = n R T

0.9934 x 0.237 = n x 0.0821 x 373

n = 7.69 x 10^-3

moles = 7.69 x 10^-3

mass / molar mass = 7.69 x 10^-3

molar mass = mass / 7.69 x 10^-3

= 0.548 / 7.69 x 10^-3

= 71.28 g/mol

molecular mass = 71.28 g / mol